OpenStudy (ttop0816):

please answer! If the legs of a right triangle are 8 and 10, find the length of the hypotenuse. the answer has to be in a square root!

hero (hero):

If you have  Use pythagorean theorem right? \[a^2 + b^2 = c^2\] \[8^2 + 10^2 = x^2\] Now solve for x: \[64 + 100 = x^2\] \[164 = x^2\] Then take square root of both sides. Don't forget to simplify.

OpenStudy (ttop0816):

thank you so much!

hero (hero):

What did you get for the hypotenuse?